Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day13


Crew 102 Log Book for April 7, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status:  The crew member with infection/inflammation problems has almost fully recovered. The rest of the Crew is doing fine.

Time Departed/Returned from EVA: No EVA, by far too windy.

Brief Narrative of Field Mission Results:

EVA Data/Interpretations:

All data collected during EVAs are being analyzed for final reports

Engineering/Hab Maintenance:

New water supply has been brought onsite by Don

Downstairs furnace still off but wall mounted backup heater is fine.

Too much water is needed to fill Green Hab aquatic tanks. The green water piping from the Hab to the gray water tank need to be restored

Plans for Tomorrow:

One EVA to unexplored cynobacteria sites

Proceed on Radiotelescope data collection

Proceed on Spirulina algae growing study

Write final report

Support Requested:

Don Lusco for:

1) Fix downstairs furnace

2) Restored the green water piping from the Hab to the gray water tank


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.58

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.47

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.54

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage): 0/4, to be refilled

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14

Trailer Potable Water Tank Level (inches from bottom): full

Hab Potable Water Tank Level (inches from bottom): 30

Potable Water Meter Reading: 44749,9

In to GreenHab Meter Reading:

Notes/Comments:

Don was back yesterday evening with the new potable water supply.


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray Water System: We started feeding the first aquatic tank but all of them are at such low level (because of evaporation) that we need plenty of gray water to fill the 3 of them. The external gray water thank is now at a level in which the manual pump do not suck anymore water. At this point we need to wait for Don to connect again the gray waters pipes of the Hub to the gray water tank.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

Boxes containing the telescope mount and controller has been handed over to Don yesterday evening


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):no

Tire Status: good

Odometer at end of day: 123903.8

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: good

Comments:


ATV 2 (Spirit):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: good

Comments:


ATV 3 (Viking I):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: good

Comments: this ATV does not start


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 76

Actual temperature upstairs (Farenheit): 75

Thermostat setting downstairs (Farenheit): N/a

Actual temperature downstairs (Farenheit): 70

Heating and Ventilation Notes/Comments:

Downstairs furnace is stopped most probably for a fan not working properly.

We are using as a backup the wall mounted heater which is working fine. At the moment the heater thermostat is in position "5"


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Science Report (Biology)

Lara Vimercati and Kavya K Manyapu Reporting

Science objective for the day was to perform a DNA extraction out of 3 samples of cyanobacteria collected

north east of the hab (UTM 12S 0520073 4251636). Each sample was collected according to sterile technique in different areas of the same site. Each sample formed a thin crust on the soil and displayed a different morphology.

Another DNA extraction was performed on Spirulina (Arthrospira platensis) samples exposed to UV light, 254 nm (1.5 ml of culture) for respectively 0,10 and 30 minutes to check on the DNA damage after exposure.

DNA extraction was performed using the MoBIO Power Soil Isolation Kit for all the samples. Verification of DNA isolation will be performed at NASA Ames research Center after the end of simulation.


Chef's Report

Franco Carbognani Reporting

1a) Today was

- a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2)  Foods served at today's dinner:

Three different kinds of pizza:

- with mushroom

- with turkey pepperoni

- with spinach/almonds pesto

no special beverage

3) List any main meal foods not finished by the crew.

Just a tiny piece of spinach/almonds pesto pizza (but probably it will not survive till midnight......)

5)  Lessons learned

You never make enough Martian Pizza.



Journalist Report

Kavya Manyapu Reporting

Windy, cold, isolated, dehydrated food, what else can you ask for on the Red Planet. The final 2 days for crew 102 are fast ending. The crew has decided to celebrate the last 2 days with a home- made (red planet) pizza with the available ‘delicious’ dehydrated-vacuum packed food. The crew is sure going to miss their commander prepared tasty Italian food.  The crew is encouraging the commander to write a book on  “how to make delicious food from crappy ingredients for isolated missions”.

Apart from the food experimentation, the crew has finally wrapped up most of the experiments. The crew biologist L.Vimercati extracted DNA from the cyanobacteria samples. K.Manyapu assisted with the DNA extraction process and learnt many things during the experiment.  Y.Brodsky has been analyzing data from the bio-erosion and corrosion samples collected. F.Carbognani has been in the process of integrating radio telescope signals into the habitat monitoring system.  A.Kapoglou has been working on a document for an outreach article in the Greek news paper. K.Manyapu has been putting together a final presentation for outreach purposes for Crew 102.

The crew has also been enjoying the last days on the red planet sparing some time for games and movies at the end of the day.

With all the constraints, environmental conditions and other tough compromises, the crew still does not regret spending these past 15 days on the red planet. The crew has thrived, survived and contributed to research for human space exploration.